# Computational Sentence-level Metrics Predicting Human Sentence Comprehension

### v1

@@ B. Attention-aware approach and its memory capability
This section provides a comprehensive guide on calculating sentence relevance using an attention-aware approach, executed in a two-step strategy. The initial step involves generating sentence embeddings using m-BERT and mGPT, followed by computing the similarity between sentences based on these embeddings. The subsequent step entails applying the attention-aware approach to manage multiple similarity values across several sentences within a window stack.
@@
In the first step, generally, using cosine similarity to compare the sentence embeddings generated by m-BERT or mPGT is a common approach for computing semantic similarity between sentences, formalized as Equation (5).
@@
$e_{s}=\text{BERT}_{\text{pooler}}(s)=\frac{1}{n}\sum_{i=1}^{n}h_{i}$ (1)
@@
After completing the first step, we elaborated on applying the attention-aware method to calculate sentence relevance. Upon acquiring the similarity values for sentences within the window stack, we proceeded to apply weights to these values. The process involves aggregating the weighted similarity values to derive a final score, as detailed in the main text. Our aim is to underscore the efficacy of the attention-aware approach in capturing contextual nuances and to explain the underlying mechanism of the weighting system.
Fig.3 shows how the attention-aware approach we adopted simulates a short-term memory stack, mirroring how readers retain memory of previously encountered words and their meanings. The method for using various weights of semantic relevance between any two sentences is inspired by both the attention mechanism found in Transformers (Panel A in Fig.3) and the human process of forgetting (Panel B in Fig. 3). The attention-aware approach can facilitate effective incorporation of contextual information and enabling it to achieve memory storage, retrieval, and integration. The weight values gradually decrease with the distance between the target sentence and the contextual sentences (see Equation (1) and the Panel B of Fig. 1), similar to the forgetting curve (Loftus, 1985).
As illustrated in the Panel B of Fig. 1, in a window stack, sentences closer to the target sentence resemble the initial days in the forgetting curve, while more distant words resemble the latter days. To simulate human forgetting mechanism, we allocated larger weights to the closer sentences and smaller weights to the distant sentences and the similarity between human forgetting mechanism and attentional weights adopted in the current study. The attention-aware approach could be linked to memory models in terms of how memory is decayed during the encoding of information, which subsequently affects how humans process sentences during reading. More importantly, the attention-aware approach is computational and fundamentally memory-based, facilitating memory storage, retrieval, and integration. This approach not only realizes the memory function but also incorporates the expectation effect.
@@
C. Statistical methods
@@
Comparing two GAMM (or LMER) models, where one is larger than the other (having all the parameters of the other model and some additional ones), the likelihood will always be higher for the larger model. This is because a larger model can fit the data better by having more parameters. However, directly comparing loglikelihoods is not appropriate when models have different sizes. Both AIC and BIC address this issue by incorporating a penalty for the number of parameters. AIC and BIC have different principles for penalizing complexity. Generally, BIC penalizes complexity more strongly than AIC, tending to favor smaller models unless both approaches agree. In essence, AIC is preferable when the main goal is prediction quality, as a slightly larger model can still provide good predictions, while a too small model usually does not. On the other hand, BIC aims to identify a reasonably sized true model by prioritizing parsimony. BIC is often better in finding the true model, but it has a higher chance of selecting a model that is too small, which is not favorable for prediction. In practice, the true model is often not “small”, but for reasons such as interpretability, smaller models are sometimes preferred even if they have slightly worse prediction performance, in which case BIC may be preferred. In summary, from the perspective of AIC, it is better to fit a slightly larger model than a too small one for improved prediction quality. However, from the perspective of BIC, both excessively large and excessively small models are equally undesirable.
@@
It is important to highlight some potential issues when using $\Delta$LogLik as a criterion. Wilcox et al. (2020) utilized GAMMs for their analysis but did not incorporate any random variables in these models. Consequently, confirming the optimality of these GAMMs becomes challenging. On the other hand, Oh and Schuler (2023) mentioned that they employed Linear Mixed Effects Models (LMER) and included random effects. However, the LMERs used in Oh and Schuler (2023) did not include “word frequency” as a control predictor. It is well-established in psycholinguistics and cognitive science that both “word length” and “word frequency” are significant variables in predicting reading time. Including these two factors as control predictors is commonly practiced when studying reading time. Therefore, when investigating reading time, it is advisable to include both “word length” and “word frequency” in GAMMs or LMERs. Meanwhile, an optimal model should also include random variables. Failure to do so may result in suboptimal GAMMs or LMERs for studying reading time. Moreover, The excessive application of heavy penalties (e.g., k in GAMMs) on the given metrics leads to overfitting in the mixed-effect models. For example, the heavy penalty on the specific metric results in the partial effect curve of this given metric becoming much steeper, and the AIC or likelihood in the model increasing remarkably. Conversely, removing such penalties eliminates these effects.
@@
Considering these factors, the GAMM models used in the present study include two control predictors: the mean word length for a sentence and the mean word length for a sentence. Additionally, random variables, participants in eye-tracking experiments, and languages, were included. We used fs (i.e.,random smooths) to adjust the trend of a numeric predictor in a nonlinear way, which includes random intercept and random slope. In other words, we can allow the metrics of interest to explore their effects on various levels of random variables comprehensively. We employed AIC to compare the performance of different GAMMs. The baseline model excludes the main predictor of interest (sentence surprisal or attention-aware sentence surprisal) but retains the other elements in the full GAMM model. After fitting the regression models, we calculated the $\Delta$AIC
values for each GAMM model by subtracting the AIC of the base GAMM model from that of a full GAMM model. A smaller AIC indicates better model performance. Similarly, a smaller $\Delta$AIC
also indicates better performance. Additioanlly, considering the best sentence surprisal and sentence relevance in each language, we used t tests to check wether $\Delta$AIC
values for two metrics have significant differences. The result shows that there is no significance difference. In other words, it is not easy to distinguish sentence surprisal or sentence relevance could predict reading speed better or not.
@@
Figure 4: The partial effects of the primary predictors—sentence surprisal and sentence relevance—on reading speed across 13 languages (i.e., Dutch, Estonian, Finnish, German, Greek, Hebrew, Italian, Korean, Norwegian, Russian, Spanish, Turkish). Note: The upper section of the diagram features sentence surprisal, while the lower portion is dedicated to sentence relevance.The x-axis signifies the computational metric, while the y-axis delineates the reading speed. To achieve a closer approximation to a normal distribution, and consequently improve the fitting, all metrics undergo a logarithmic transformation. Each curve visually articulates the correlation between a predictor variable and the response variable, namely reading speed. A steeper incline on these curves underscores a more robust impact between the predictor and reading speed, whereas gentler slopes imply a less pronounced effect. Moreover, when a curve fluctuates around zero, its effect vanishes. The information regarding $\rho$-values and $\Delta$AIC is displayed at the top of each plot. The methodology for calculating $\Delta$AIC for “sentence surprisal” and “sentence relevance” is detailed in the main text. In conclusion, sentence surprisal seems to lack significance in English, Korean, and Russian. However, sentence relevance may show no significant impact in English, Russian, Spanish, and Turkish.
D. Correlation between sentence surprisal and sentence relevance

### v2

@@ D. Attention-aware approach and its memory capability
The initial step involves generating sentence embeddings using m-BERT or mGPT, followed by computing the similarity between two sentences based on these embeddings. The subsequent step entails applying the “attention-aware” approach to manage multiple similarity values across several sentences within a window stack.
@@
In the first step, we need to generate embeddings for each sentence in this window by employing BERT or GPT. After obtaining sentence embedding for each sentence, we applied cosine similarity to compare the sentence embeddings generated by m-BERT or mPGT, and this practice is a common approach for computing semantic similarity between sentences, formalized as Equation (7).
@@
$e_{s}=\text{BERT}_{\text{pooler}}(s)=\frac{1}{n}\sum_{i=1}^{n}h_{i}\quad(8)$
@@
The similarity score between any two sentences in this window is calculated. Following this, four cosine similarity values are obtained for the sentences. Subsequently, we implemented the second step.
After completing the first step, we elaborated on applying weights to calculate sentence relevance. Upon acquiring the similarity values for sentences within the window stack, we proceeded to apply weights to these values. The process involves aggregating the weighted similarity values to derive a final score, as detailed in the main text. Our aim is to underscore the efficacy of the attention-aware approach in capturing contextual information and to explain the underlying mechanism of the weighting system.
As illustrated in the Panel B of Fig. 1, in a window stack, sentences closer to the target sentence resemble the initial days in the forgetting curve, while more distant words resemble the latter days. To simulate human forgetting mechanism, we allocated larger weights to the closer sentences and smaller weights to the distant sentences and the similarity between human forgetting mechanism and attentional weights adopted in the current study. The weight values gradually decrease with the distance between the target sentence and the contextual sentences (see Equation (1) and the Panel B of Fig. 1), similar to the forgetting curve (Loftus, 1985). The attention-aware approach could be linked to memory models in terms of how memory is decayed during the encoding of information, which subsequently affects how humans process sentences during reading. More importantly, the attention-aware approach is computational and fundamentally memory-based, facilitating memory storage, retrieval, and integration. This approach not only realizes the memory function but also incorporates the expectation effect.
Furthermore, when applying a 3-sentence window (the target sentence is not takne into account), akin to the fading of memories over three days, the average rate of memory decline can be estimated to average one-third per day. This fading initiates more abruptly and slows as time progresses. By modeling this, we can recalibrate the significance of sentences based on their distance from the target sentence, diminishing their value by a one-third. In this sense, sentences further from the target sentence receive lower weights compared to those in a uniform distribution. This leads to a gradual weighting scale, such as 1/3, 2/3, 1/2, and so forth, mirroring their relative proximity to the target sentence.
The Panel A of Fig.4 shows how the attention-aware approach we adopted simulates a short-term memory stack, mirroring how readers retain memory of previously encountered words and their meanings. The method for using various weights of semantic relevance between any two sentences is inspired by both the attention mechanism found in Transformers (Panel A in Fig.4) and the human process of forgetting (Panel B in Fig. 4). The attention-aware approach can facilitate effective incorporation of contextual information and enabling it to achieve memory storage, retrieval, and integration.
Finally, as previously noted, the term “attention-aware” bears a resemblance to the algorithm of attention in Transformers, primarily because the attention-aware approach we proposed significantly enhances computational efficiency through the incorporation of contextual information. However, our attention-aware method does not incorporate any attention layers from Transformers. Each step in the computation of attention-aware metrics is transparent and interpretable.
@@
E. Statistical methods and comparison standards
@@
Comparing two GAMM (or LMER, Linear Mixed Effects Models) models, where one is larger than the other (having all the parameters of the other model and some additional ones), the likelihood will always be higher for the larger model. This is because a larger model can fit the data better by having more parameters. However, directly comparing loglikelihoods is not appropriate when models have different sizes. Both AIC and BIC address this issue by incorporating a penalty for the number of parameters. AIC and BIC have different principles for penalizing complexity. Generally, BIC penalizes complexity more strongly than AIC, tending to favor smaller models unless both approaches agree. In essence, AIC is preferable when the main goal is prediction quality, as a slightly larger model can still provide good predictions, while a too small model usually does not. On the other hand, BIC aims to identify a reasonably sized true model by prioritizing parsimony. BIC is often better in finding the true model, but it has a higher chance of selecting a model that is too small, which is not favorable for prediction. In practice, the true model is often not “small”, but for reasons such as interpretability, smaller models are sometimes preferred even if they have slightly worse prediction performance, in which case BIC may be preferred. In summary, from the perspective of AIC, it is better to fit a slightly larger model than a too small one for improved prediction quality. However, from the perspective of BIC, both excessively large and excessively small models are equally undesirable (Vrieze, 2012).
Moreover, the developer of R package on GAMM (“mgcv”) used AIC to make model comparison (Wood et al., 2016; Wood, 2020). AIC has also been mostly taken to understand model performance in psycholinguistic research (Wieling, 2018; Baayen and Linke, 2020 and the relevant studies) if GAMM or generalized mixed-effect models are employed.
@@
We should point out some potential issues when using $\Delta$LogLik as a criterion. Wilcox et al. (2020) utilized GAMMs for their analysis but did not incorporate any random variables in these models. Consequently, confirming the optimality of these GAMMs becomes challenging. On the other hand, Oh and Schuler (2023) mentioned that they employed LMER and included random effects. However, the LMERs used in Oh and Schuler (2023) did not include “word frequency” as a control predictor. It is well-established in psycholinguistics and cognitive science that both “word length” and “word frequency” are significant variables in predicting reading time. Including these two factors as control predictors is commonly practiced when studying reading time. Therefore, when investigating reading time, it is advisable to include both “word length” and “word frequency” in GAMMs or LMERs. Meanwhile, an optimal model should also include random variables. Failure to do so may result in suboptimal GAMMs or LMERs for studying reading time. Moreover, The excessive application of heavy penalties (e.g., k in GAMMs) on the given metrics leads to overfitting in the mixed-effect models. For example, the heavy penalty on the specific metric results in the partial effect curve of this given metric becoming much steeper, and the AIC or likelihood in the model increasing remarkably. Conversely, removing such penalties eliminates these effects.
@@
Considering these factors, the GAMM models used in the present study include two control predictors: the mean word length for a sentence and the mean word length for a sentence. Additionally, random variables, participants in eye-tracking experiments, and languages, were included. We used fs (i.e.,random smooths) to adjust the trend of a numeric predictor in a nonlinear way, which includes random intercept and random slope. In other words, we can allow the metrics of interest to explore their effects on various levels of random variables comprehensively. We employed AIC to compare the performance of different GAMMs. The baseline model excludes the main predictor of interest (sentence surprisal or attention-aware sentence surprisal) but retains the other elements in the full GAMM model. After fitting the regression models, we calculated the $\Delta$AIC values for each GAMM model by subtracting the AIC of the base GAMM model from that of a full GAMM model. A smaller AIC indicates better model performance. Similarly, a smaller $\Delta$AIC also indicates better performance. Additioanlly, considering the best sentence surprisal and sentence relevance in each language, we used t tests to check wether $\Delta$AIC values for two metrics have significant differences. The result shows that there is no significance difference. In other words, it is not easy to distinguish sentence surprisal or sentence relevance could predict reading speed better or not.
F. Performance of individual languages
The performance of an individual language is shown in Fig. 5. There are two key standards for interpreting the curve of partial effect in each plot. The first standard involves analyzing the curve’s steepness. A steeper incline signifies a stronger correlation between the predictor and reading speed, whereas a gentler slope indicates a less pronounced effect. The second standard focuses on the curve’s fluctuation around zero; a curve that hovers around zero suggests its impact is minimal. For instance, in the English plot, a curve consistently near zero would indicate a weak effect. Understanding mixed-effect models is essential for appreciating the significance of these indicators, highlighting the Akaike Information Criterion (AIC) as the preferred tool for model comparison.
As shown in Fig. 5, it appears that sentence surprisal lacks significance in English, Korean, and Russian, while sentence relevance shows no significant impact in English, Russian, Spanish, and Turkish.
@@ F. Performance of individual languages
Figure 5: The partial effects of the primary predictors—sentence surprisal and sentence relevance—on reading speed across 13 languages (i.e., Dutch, Estonian, Finnish, German, Greek, Hebrew, Italian, Korean, Norwegian, Russian, Spanish, Turkish). Note: The upper section of the diagram features sentence surprisal, while the lower portion is dedicated to sentence relevance.The x-axis signifies the computational metric, while the y-axis delineates the reading speed. To achieve a closer approximation to a normal distribution, and consequently improve the fitting, all metrics undergo a logarithmic transformation. Each curve visually articulates the correlation between a predictor variable and the response variable, namely reading speed. A steeper incline on these curves underscores a more robust impact between the predictor and reading speed, whereas gentler slopes imply a less pronounced effect. Moreover, when a curve fluctuates around zero, its effect vanishes. The information regarding p-values and $\Delta$AIC is displayed at the top of each plot. The methodology for calculating $\Delta$AIC for “sentence surprisal” and “sentence relevance” is detailed in the main text. In conclusion, sentence surprisal seems to lack significance in English, Korean, and Russian. However, sentence relevance may show no significant impact in English, Russian, Spanish, and Turkish.
Next, we employed T-test to assess the statistical significance of the $\Delta$AIC values for both sentence surprisal and sentence relevance. The analyses revealed no significant differences, with the p-value substantially exceeding 0.05. This outcome suggests that sentence surprisal and sentence relevance may contribute equivalently to the prediction of reading speeds. As a result, it can be inferred that both metrics are effective in facilitating the overall comprehension of sentences. Despite this, Table 2 shows that the overall predictive power of sentence surprisal with m-BERT and CR is stronger than sentence relevance.
G. Correlation between sentence surprisal and sentence relevance